# Byung-Kwan Lee KAIST

### v2

@@ 4 Experiment
We have demonstrated super vision language performances of Refer to captionTroL, despite its limited model size, as depicted in Table 1, Table 2, and Table 4. This is attributed to the enhanced learning capabilities by layer traversing. Furthermore, Table 4 have shown several ablation studies to clearly corroborate the effectiveness of Refer to captionTroL in light of seven factors: (a) backbone pre-trained LLMs, (b) the use of layer traversing, (c) the structure of TroL Gating, (d) mixing operation of TroL-Mixer, (e) training percentage of visual instruction tuning, and (f) ratio of measured time for inference speed. Note that, in Table 4(c), ‘FC’ denotes a fully-connected layer, and ‘MHA’ denotes a multi-head attention block, and ‘PR’ denotes the structure of Perceiver Resampler (Alayrac et al., 2022). Through the ablation, we consider the efficiency of the model structure and performance on vision language tasks to build the current architectures of Refer to captionTroL, including the pre-trained LLM, TroL Gating architecture, and the TroL-Mixer operation. Additionally, layer traversing focusing on user’s questions during inference makes text generation speed comparable to the baselines we used for pre-trained LLM. Note that, all descriptions of the evaluation benchmarks used in this paper are explained in Appendix A, and we show diverse samples for Refer to captionTroL’s text generation quality in Appendix B.
@@ 5 Discussion and Conclusion
A new LLVM family, Refer to captionTroL, has demonstrated significant advancements in vision language performance despite its inherent limitation of having smaller layers compared to larger model sizes in both open- and closed-source LLVMs. Table 4 suggests that reusing layers through layer traversing can be an effective alternative to incorporating additional modules, such as mixed vision encoders and multiple computer vision models. We expect Refer to captionTroL to remain a competitive and efficient option in the rapidly evolving field of vision language models, solidifying its place in the landscape of efficient LLVMs.

### v3

@@ 4 Experiment
We have demonstrated super vision language performances of Refer to captionTroL, despite its limited model size, as depicted in Table 1, Table 2, and Table 4. This is attributed to the enhanced learning capabilities by layer traversing. Furthermore, Table 4 have shown several ablation studies to clearly corroborate the effectiveness of Refer to captionTroL in light of seven factors: (a) backbone pre-trained LLMs, (b) the use of layer traversing, (c) the structure of TroL Gating, (d) mixing operation of TroL-Mixer, (e) training percentage of visual instruction tuning, and (f) ratio of measured time for inference speed. Note that, in Table 4(c), ‘FC’ denotes a fully-connected layer, and ‘MHA’ denotes a multi-head attention block, and ‘PR’ denotes the structure of Perceiver Resampler (Alayrac et al., 2022). Through the ablation, we consider the efficiency of the model structure and performance on vision language tasks to build the current architectures of Refer to captionTroL, including the pre-trained LLM, TroL Gating architecture, and the TroL-Mixer operation. Additionally, layer traversing focusing on user’s questions during inference makes text generation speed comparable to the baselines we used for pre-trained LLM. Note that, all descriptions of the evaluation benchmarks used in this paper are explained in Appendix A, and we show diverse samples for Refer to captionTroL’s text generation quality in Appendix B. In addition, Appendix C provides further ablation studies to check more various factors.
@@ 5 Discussion and Conclusion
A new LLVM family, Refer to captionTroL, has demonstrated significant advancements in vision language performance despite its inherent limitation of having smaller layers compared to larger model sizes in both open- and closed-source LLVMs. Table 4 suggests that reusing layers through layer traversing can be an effective alternative to incorporating additional modules. We expect Refer to captionTroL to remain an efficient option in the rapidly evolving field, solidifying its place in the landscape of efficient LLVMs.
@@
Acknowledgments
This work was partially supported by two funds: Center for Applied Research in Artificial Intelligence (CARAI) grant funded by DAPA and ADD (UD230017TD) and IITP grant funded by the Korea government (MSIT) (RS-2022-II220984). Additionally, it was supported by the KISTI National Supercomputing Center with supercomputing resources including technical support (KSC-2024-CRE-0160).
@@ Appendix B Refer to captionTroL Generation Quality
Appendix C Further Ablation Studies
[…]
TroL-1.8B w.o. Lay-Trav 0 36.0 34.6 75.0
TroL-1.8B w. MoE (4 Experts, Top k=3) 402M 39.1 38.8 80.2
TroL-1.8B w. Lay-Trav 49K 45.5 45.1 87.5
TroL-3.8B w.o. Lay-Trav 0 37.4 35.3 81.4
TroL-3.8B w. MoE (4 Experts, Top k=3) 1.2B 39.9 41.1 83.4
TroL-3.8B w. Lay-Trav 98K 46.5 51.0 90.8
TroL-7B w.o. Lay-Trav 0 41.2 45.8 82.7
TroL-7B w. MoE (4 Experts, Top k=3) 2.1B 45.6 50.5 86.6
TroL-7B w. Lay-Trav 131K 51.3 54.7 92.8
Table 5: Performance comparison of Refer to captionTroL across various benchmarks.
Inspired by the Mixture of Experts (MoE) concept (Shazeer et al., 2017; Lin et al., 2024), we designed the architecture of TroL-Gating and TroL-Mixer. The primary advantage of using MoE is achieving significant performance improvements without adding many physical layers or significantly increasing model sizes.
The paragraph below Equation 2 in Shazeer et al. (2017) states: ‘A simple choice of non-sparse gating function is to multiply the input by a trainable weight matrix and then apply the Softmax function.’ Here, the concept of a ‘gating function’ is analogous to TroL-Gating. Similarly, Equation 6 in Lin et al. (2024) performs a weighted average on the features obtained from each expert module, which is akin to the weighted average operation in TroL-Mixer.
The only difference between traditional MoE and TroL is whether multiple expert modules are physically used. In TroL, the first propagation and the second propagation are implicitly considered as two expert modules in MoE. We believe this iterative propagation introduces a broader range of vision-language knowledge, as layer traversing handles more parameters than MoE. Table 5 below compares the MoE and layer traversing techniques.
Therefore, we conclude that the layer traversing technique embeds more vision-language knowledge than MoE without adding many physical layers or significantly increasing model sizes. This explains why layer traversing performs well across general tasks. Moreover, this approach intuitively matches the human process of retracing answering stream. We hope the layer traversing technique will be regarded as a promising direction for future MoE research.
[…]
DeepSeek-VL-1.3B 39.9 34.8 51.1
DeepSeek-VL-1.3B 45.9 46.2 60.5
MiniCPM-2.8B 39.1 41.0 69.2
MiniCPM-2.8B 47.4 50.2 80.8
LLaVA-NeXT-7B 40.2 43.9 72.3
LLaVA-NeXT-7B 49.8 52.7 84.4
Table 6: Performance comparison of various models with and without Lay-Trav across benchmarks.
In Table 6, we have applied layer traversing to other LLVMs such as DeepSeek-VL-1.3B, MiniCPM-V2-2.8B, and LLaVA-NeXT-7B. This adjustment aims to validate the effectiveness of layer traversing under a fairer comparison setting, using the same TroL visual instruction tuning dataset.
[…]
TroL-1.8B (Backbone Model) 25.1 21.4 44.6
TroL-1.8B (Backbone Model) 24.8 15.9 36.0
TroL-1.8B 36.0 34.6 75.0
TroL-1.8B 45.5 45.1 87.5
TroL-3.8B (Backbone Model) 26.6 22.3 45.4
TroL-3.8B (Backbone Model) 25.2 16.0 36.1
TroL-3.8B 37.4 43.5 78.4
TroL-3.8B 46.5 51.1 90.8
TroL-7B (Backbone Model) 27.3 21.1 51.3
TroL-7B (Backbone Model) 24.1 15.3 37.1
TroL-7B 41.2 45.8 82.7
TroL-7B 51.3 54.7 92.8
Table 7: Performance comparison of Refer to captionTroL with and without Lay-Trav and different training setups.
In Table 7, we evaluated the experiments you asked in the table below. Without any training, layer traversing technique confuses LLM performances because it never see this kind of features.
[…]
TroL-1.8B 2 45.5 45.1 87.5
TroL-1.8B 3 45.7 45.6 87.8
TroL-1.8B 4 45.8 45.9 88.0
TroL-3.8B 2 46.5 51.1 90.8
TroL-3.8B 3 46.6 51.4 91.1
TroL-3.8B 4 46.9 51.8 91.2
TroL-7B 2 51.3 54.7 92.8
TroL-7B 3 51.7 55.0 93.2
TroL-7B 4 52.0 55.2 93.3
Table 8: Performance comparison of Refer to captionTroL across different numbers of propositions.
The reason we use the second propagation instead of the third propagation is that marginal improvements are observed when using more than two propagation. Table 8 shows the performance across the propagation numbers from 2 to 4.
[…]
TroL-1.8B (Question) 1.0 45.5 45.1 87.5
TroL-1.8B (Question-Answer) 4.9 46.3 47.2 89.2
TroL-3.8B (Question) 1.0 46.5 51.1 90.8
TroL-3.8B (Question-Answer) 5.1 47.9 53.4 92.1
TroL-7B (Question) 1.0 51.3 54.7 92.8
TroL-7B (Question-Answer) 5.7 52.5 56.2 94.3
Table 9: Performance comparison of Refer to captionTroL across question and question-answer setups.
The primary reason for not applying layer traversing to the answer part is the huge increase in answering time complexity. Fortunately, turning layer traversing on or off for the answer part during inference shows a little significant performance gap in Table 9, therefore it is better for us to deal with only question part in efficient inference.
[…]
TroL-1.8B 36.0 34.6 75.0
TroL-1.8B (Nothing) 42.1 41.0 82.8
TroL-1.8B (Question) 45.5 45.1 87.5
TroL-1.8B (Question-Answer) 46.3 47.2 89.2
TroL-3.8B 37.4 43.5 78.4
TroL-3.8B (Nothing) 43.8 46.3 86.3
TroL-3.8B (Question) 46.5 51.1 90.8
TroL-3.8B (Question-Answer) 47.9 53.4 92.1
TroL-7B 41.2 45.8 82.7
TroL-7B (Nothing) 47.0 50.8 88.9
TroL-7B (Question) 51.3 54.7 92.8
TroL-7B (Question-Answer) 52.5 56.2 94.3
Table 10: Performance comparison of Refer to captionTroL with Lay-Trav in training and various configurations.
In Table 10, thanks to layer traversing together with visual instruction tuning, we observed that vision language performance improves even when question and answer traversing are turned off.